### v2

# Flyby-induced displacement: analytic solution

Abstract
Approximating the derivative-of-the-Gaussian profile proposed by Gibbons and Hawking by the hyperbolic Scarf potential, the scattering of particles by a gravitational wave generated by flyby is described analytically by following the Nikiforov-Uvarov method applied to the Scarf potential. Pure displacement arises when the wave zone contains an integer number of half-waves. The results confirm the prediction of Zel’dovich and Polnarev.
Key words: gravitational waves, displacement memory effect, flyby, Nikiforov-Uvarov method
pacs:
04.20.-q Classical general relativity;
preprint: arXiv: 2502.0136v2 [gr-qc]
I Analytic solution for flyby
Zel’dovich and Polnarev suggested that gravitational waves generated by flyby could be observed through the displacement memory effect (DM) ZelPol ; DM-1 ; Jibril . Since then, and with the discovery that gravitational memory effects are closely related to soft graviton theorems and Bondi-Metzner-Sachs symmetries at future null infinity, the field has grown considerably BMS1 ; BMS2 ; BMS3 ; BMS4 . Zel’dovich - Polnarev suggestion was confirmed numerically for the derived-of-the-Gaussian profile ${\cal A}^{dG}\propto d\big{(}\exp[-U^{2}]\big{)}/dU$ proposed by Gibbons and Hawking GibbHaw71 provided the parameters take some “magic” values DM-1 . Their profile does not support an analytic solution, however it can be approximated by the derived Pöschl-Teller potential PTeller ; Chakra ; DM-1 ,
${\cal A}^{dPT}(U)=\frac{\;d}{dU}\left(\frac{g}{2\cosh^{2}U}\right)=-g\frac{\sinh{U}}{\cosh^{3}{U}}\,,$ (1)
which can be studied semi-analytically using confluent Heun functions DM-1 ; DM-2 . This led us to conjecture a sort of quantisation condition: DM ariseses when the wave zone contains an integer number $m$ of half-waves. In this note we present yet another approximation of the flyby profile, namely by the hyperbolic Scarf potential scarfpot ,
${\cal A}^{scarf}(U)=-2g\frac{\sinh{U}}{\cosh^{2}{U}}\,,$ (2)
for which an analytic solution is found; it confirms the above conjecture.
For small values of $U$ it behaves as (1) but for large $U$ it falls off more slowly. However dilating $U$ appropriately almost perfect overlapping is obtained, as depicted in FIG.1.
Figure 1: The derived Pöschl-Teller  profile (1) and its Scarf potential alter-ego (2) are both good approximations of the derived-Gaussian one proposed by Gibbons and Hawking GibbHaw71 to describe flyby.
In detail, the geodesics are, in Brinkmann coordinates, determined by the a Sturm-Liouville-type transverse equations DM-2 ,
$\dfrac{d^{2}\!X^{i}}{dU^{2}}-\frac{1}{2}{\cal A}X^{i}=0\,,$ (3)
$i=1,2$, where $U$ behaves as non-relativistic time and ${\cal A}={\cal A}(U)$ is the profile of the gravitational wave DBKP ; DGH91 . These equations are completed with the DM boundary conditions,
$\frac{d{\bm{X}}}{dU}\Big{|}(U=-\infty)=0=\frac{d{\bm{X}}}{dU}\Big{|}(U=+\infty)\,.$ (4)
Below we show that the hyperbolic Scarf potential scarfpot ; Scrf1 , eqn. (2), allows us to find an analytic solution which confirms the quantisation conjecture above. Eqn.(3) can be viewed indeed as a Schrödinger equation for the (non-normalizable) zero-energy ground state DM-1 . The solution is obtained by switching to $t=\sinh{U}$ which takes (3) into  [^4] For comparison, putting $t=\tanh U$ carries (3) with Pöschl-Teller profile ${\cal A}^{PT}=(g/2)\cosh^{-2}U$ into a form which can be solved by Legendre polynomials DM-1 .
$(1+t^{2})\frac{d^{2}X}{dt^{2}}+t\frac{dX}{dt}-\frac{gt}{1+t^{2}}X=0\,,$ (5)
which is of the generalised hypergeometric type, and can be solved analytically by the Nikiforov-Uvarov method NikiUvar . Since a fairly detailed description of this simple, yet elegant and powerful method can be found elsewhere NikiUvar ; NU1 ; NU2 ; NU3 ; NU4 , we only provide a brief summary of the algorithm in the Appendix.
The Nikoforov-Uvarov method (recalled in the Appendix) applies to second-order differential equations of the generalised hypergeometric type of the form,
$u^{\prime\prime}+\frac{\pi_{1}(z)}{\sigma(z)}u^{\prime}+\frac{\sigma_{1}(z)}{\sigma^{2}(z)}u=0\,.$ (6)
For Eq. (5) in particular,
$\sigma=1+t^{2},\;\;\;\pi_{1}=t,\;\;\;\sigma_{1}=-gt\,.$ (7)
Then,
$\sigma_{3}=\frac{t^{2}}{4}+gt+k(1+t^{2})=t^{2}\left(k+\frac{1}{4}\right)+gt+k\,,$ (8)
Therefore, the condition $\det(\sigma_{3})=g^{2}-4k\left(k+\frac{1}{4}\right)=0$ gives
$k=\frac{\pm\sqrt{1+16g^{2}}-1}{8},$ (9)
and
$|g|=2\sqrt{k\left(k+\frac{1}{4}\right)}\,.$ (10)
Thus we have
$\sigma_{3}=\frac{g^{2}}{4k}\left[t^{2}+\frac{4k}{g}t+\frac{4k^{2}}{g^{2}}\right]=\frac{g^{2}}{4k}\left(t+\frac{2k}{g}\right)^{2}\,\quad\pi=\frac{t}{2}\pm\sqrt{\sigma_{3}}=\frac{t}{2}\pm\frac{|g|}{2\sqrt{k}}\left|t+\frac{2k}{g}\right|\,.$ (11)
Then
$\tau=\pi_{1}+2\pi=2t\pm\frac{|g|}{\sqrt{k}}\left|t+\frac{2k}{g}\right|.$ (12)
To have real solutions, we assume $k>0$, that is we choose upper sign in Eq.(9): $k=\frac{\sqrt{1+16g^{2}}-1}{8}.$ On the other side, the quantization condition Eq.(34) requires
$\lambda=-n(n+1)\mp\epsilon\,\frac{n|g|}{\sqrt{k}},$ (13)
since $\tau^{\prime}=2\pm\epsilon\,\frac{|g|}{\sqrt{k}}$ with $\epsilon=\mathrm{sign}\left(t+\frac{2k}{g}\right)\,.$ Therefore we get, after some algebra,
$k=\lambda-\pi^{\prime}=-n^{2}-n-\frac{1}{2}\mp\epsilon\frac{|g|}{\sqrt{k}}\left(n+\frac{1}{2}\right),$
and using Eq. (10) for $|g|/\sqrt{k}$, we get
$\sqrt{k+\frac{1}{4}}=\mp\epsilon\left(n+\frac{1}{2}\right)$ (14)
and we see that we must choose the lower signs in Eq.(11) and Eq.(12) to get $\mp\epsilon=1$. Then the solution of Eq.(14) is
$k_{n}=n(n+1),$ (15)
and Eq.(10) determines the “magic” strengths of the Scarf potential Eq.(2), for which we have polynomial solutions for $y_{n}(t)$ labeled by the integer “quantum” number $n$,
$\;\;\displaystyle{|g_{n}|=(2n+1)\sqrt{n(n+1)}\;.}\;\;$ (16)
It remains to find the gauge function $\varphi(t)$ and the corresponding hypergeometric type polynomials $y_{n}$. Because of Eq. (15) and Eq. (16), we have,
$\pi(t)=-nt-\mathrm{sign}(g)\sqrt{n(n+1)},\;\;\;\;\tau(t)=-(2n-1)t-2\mathrm{sign}(g)\sqrt{n(n+1)}\,.$ (17)
The equation for the gauge function,
$\frac{d\varphi}{dt}=-\frac{nt+\mathrm{sign}(g)\sqrt{n(n+1)}}{1+t^{2}}\,,$ (18)
can be easily integrated and we get (up to irrelevant additive constant which translates to a multiplicative constant in the solution $X(t)$),
$\varphi(t)=-\frac{n}{2}\ln(1+t^{2})+\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}\,.$ (19)
Therefore, the corresponding geodesic trajectory is given by
$\;\;\displaystyle{X_{n}(t)=(1+t^{2})^{-\frac{n}{2}}\,e^{-\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,y_{n}(t)\,.}\;\;$ (20)
The polynomial $y_{n}(t)$ is given by the Rodrigues formula Eq. (35) in which we need to find the weight function $\rho(t)$, for which we found,
$\rho(t)=(1+t^{2})^{-\left(n+\frac{1}{2}\right)}e^{-2\mathrm{sign}(g)\,\sqrt{n(n+1)}\arctan{t}}\,.$ (21)
Therefore
$\;\;\displaystyle{y_{n}(t)=\frac{B_{n}}{\rho(t)}\left[(1+t^{2})^{n}\rho(t)\right]^{(n)}\,.}\;\;$ (22)
The normalization constants $B_{n}$ should be determined from the initial conditions. The first three polynomials are $y_{1}(t)\sim\tau(t)=-t-2\sqrt{2}\,\mathrm{sign}(g),\;y_{2}(t)\sim 2t^{2}+8\sqrt{6}\,\mathrm{sign}(g)t+23,\;y_{3}(t)\sim-6t^{3}-72\sqrt{3}\,\mathrm{sign}(g)t^{2}-423t-172\sqrt{3}\,\mathrm{sign}(g)\,.$ They can be expressed in terms of Routh-Romanovski polynomials scarfpot ; Raposo as $y_{n}(t)=R_{n}^{(\alpha,\beta)}(t)$ with $\alpha=-2\,\mathrm{sign}(g)\sqrt{n(n+1)}$ and $\beta=-n+\frac{1}{2}$.
For (half)wave numbers $n=1,2,3,4$ the analytic solutions (20) reproduce almost identically those half-numerical ones shown in FIG.s # 8 & 9 of DM-2 for the derived Pöschl-Teller case. For $n\geq 5$, though, DM requires to find a high number of decimals and the numerical procedure becomes tedious. Plotting instead the analytic solution (20)-(22) yields FIG.2.
Figure 2: DM geodesics for the Scarf profile (2) for wave numbers $n=5$ and $n=6$.
In conclusion, our results show that the profile proposed by Hawking and Gibbons for flyby GibbHaw71 can be approximated by the Scarf potential (2) in FIG.1 which supports analytic solutions, (20)-(22), labeled by an integer $n$, (16), which can be constructed by the Nikiforov-Uvarov method. The displacement effect (DM) arises when the wave zone contains $n$ half-waves as illustrated in FIG.2, extending an observation made DM-2 for the Gaussian/Pöschl-Teller profile.
Further physical applications of the Scarf potential which include supersymmetry scarfpot ; Compean , non-central potentials Dutt and even quarks (QCD), are reviewed in ref. Raposo . Note that, in the supersymmetry literature, the Scarf hyperbolic potential has recently attracted much attention Scrf1 . This potential was discovered by Gendenshtein and is therefore referred also to as the Gendenstein potential Gendenon ; Natanson . See also Hait ; Bhattacharya23 ; Bhattacharya:2025ljc for related work.
For isotropic profile we have a “universal model” : every plane gravitational wave can be mapped locally conformally to free Minkowski space by following Niederer, or Arnold Niederer ; Arnold ; Takagi , consistently with the vanishing of the Cotton or Weyl tensor Inomata ; Dhasmana ; GibbonsDark ; ZZH21 .
However gravitational waves are isotropic only in 1d, and their anisotropy in d=2 dim implies that we arrive to unchartered territory Lukash ; Inomata ; ZZH21 and it is not clear whether we still have a (conformal) “universal ancestor”.
Acknowledgements.
The authors are grateful to Gary Gibbons for discussions. PMZ was partially supported by the National Natural Science Foundation of China (Grant No. 11975320).
Appendix A Compendium on the Nikoforov-Uvarov method
As it was already mentioned in the text, the Nikoforov-Uvarov method applies to second-order differential equations of the generalised hypergeometric type of the form
$u^{\prime\prime}+\frac{\pi_{1}(z)}{\sigma(z)}u^{\prime}+\frac{\sigma_{1}(z)}{\sigma^{2}(z)}u=0\,,$ (23)
where the prime denotes $d/dz$. $z$ can be complex. $\pi_{1}(z)$ is a polynomial at most of the first degree, and $\sigma(z)$, $\sigma_{1}(z)$ are polynomials at most of the second degree NikiUvar .
The set of solutions of Eq. (23) is invariant under “gauge” transformations $u(z)\to y(z)$, $u(z)=e^{\varphi(z)}y(z),$ if the gauge function $\varphi(z)$ satisfies the equation
$\varphi^{\prime}(z)=\frac{\pi(z)}{\sigma(z)}\,$ (24)
where $\pi(z)$ is a polynomial at most of the first degree. In this case for $y(z)$ we obtain also a generalised hypergeometric-type equation:
$y^{\prime\prime}+\frac{\pi_{2}(z)}{\sigma(z)}y^{\prime}+\frac{\sigma_{2}(z)}{\sigma^{2}(z)}y=0\,,$ (25)
where $\pi_{2}(z)=\pi_{1}(z)+2\pi(z)$ is a polynomial of at most of the first degree, and
$\sigma_{2}(z)=\sigma_{1}(z)+\pi^{2}(z)+\pi(z)\left[\pi_{1}(z)-\sigma^{\prime}(z)\right]+\pi^{\prime}(z)\sigma(z)$ (26)
is a polynomial of at most of the second degree. Choosing $\pi(z)$ so that
$\sigma_{2}(z)=\lambda\sigma(z)\,$ (27)
where $\lambda$ is a constant, Eq. (6) is simplified to an hypergeometric type equation,
$\sigma(z)y^{\prime\prime}+\pi_{2}(z)y^{\prime}+\lambda y=0\,.$ (28)
In the light of Eq. (26), Eq. (27) means
$\pi^{2}+\pi[\pi_{1}-\sigma^{\prime}]+\sigma_{1}-k\sigma=0,$ (29)
where $k=\lambda-\pi^{\prime}$ is another constant. Eq. (29) is solved by,
$\pi=\frac{\sigma^{\prime}-\pi_{1}}{2}\pm\sqrt{\sigma_{3}(z)}{\quad\text{\small where}\quad}\sigma_{3}(z)=\left(\frac{\sigma^{\prime}-\pi_{1}}{2}\right)^{2}-\sigma_{1}+k\sigma\,.$ (30)
$\pi$ is a polynomial only when $\sigma_{3}(z)$ is the square of a first order polynomial. Hence it has a double root and its discriminant is zero, $\Delta(\sigma_{3})=0\,,$ which determines the constant $k$, and hence the constant $\lambda$.
We want to find polynomial solutions of Eq. (28). It can be shown that derivatives of $y(z)$, $v_{n}(z)=y^{(n)}(z)$, are also generalised hypergeometric-type functions,
$\sigma v_{n}^{\prime\prime}+\tau_{n}(z)v_{n}^{\prime}+\mu_{n}v_{n}=0\,,$ (31)
and we get the recurrence relations
$\tau_{n}(z)=\sigma^{\prime}(z)+\tau_{n-1}(z),\;\;\;\;\mu_{n}=\mu_{n-1}+\tau_{n-1}^{\prime}\,,$ (32)
with initial values $\tau(z)=\tau_{0}(z)=\pi_{2}(z),\;\mu_{0}=\lambda\,.$ If $y(z)=y_{n}(z)$ is a polynomial of order $n$, then $v_{n}=\mathrm{const}$ and Eq. (31) is satisfied only if $\mu_{n}=0$. But the repeated application of the recurrence relations Eq. (32) produces
$\tau_{n}(z)=n\sigma^{\prime}(z)+\tau(z),\;\;\;\mu_{n}=\lambda+n\tau^{\prime}+\frac{1}{2}n(n-1)\sigma^{\prime\prime}\,.$ (33)
Getting a polynomial solution $y(z)=y_{n}(z)$ of Eq. (28) thus requires the “quantization condition”,
$\lambda=\lambda_{n}=-n\tau^{\prime}-\frac{1}{2}n(n-1)\sigma^{\prime\prime}.$ (34)
It can be shown NikiUvar that, when the quantization condition Eq. (34) is satisfied, the corresponding polynomial solution of Eq. (28) is given by the Rodrigues formula
$y_{n}(z)=\frac{B_{n}}{\rho(z)}\left[\sigma^{n}(z)\rho(z)\right]^{(n)},$ (35)
where $B_{n}$ is some (normalization) constant, and the weight functions $\rho(z)$ satisfies Pearson’s equation Ismail
$(\sigma\rho)^{\prime}=\rho\tau\,.$ (36)